# Bridging the Gap: Towards an Expanded Toolkit for ML-Supported Decision-Making in the Public Sector

### v2

@@ Abstract
Machine Learning (ML) systems are becoming instrumental in the public sector, with applications spanning areas like criminal justice, social welfare, financial fraud detection, and public health. While these systems offer great potential benefits to institutional decision-making processes, such as improved efficiency and reliability, they still face the challenge of aligning nuanced policy objectives with the precise formalization requirements necessitated by ML models. In this paper, we aim to bridge the gap between ML model requirements and public sector decision-making by presenting a comprehensive overview of key technical challenges where disjunctions between policy goals and ML models commonly arise. We concentrate on pivotal points of the ML pipeline that connect the model to its operational environment, discussing the significance of representative training data and highlighting the importance of a model setup that facilitates effective decision-making. Additionally, we link these challenges with emerging methodological advancements, encompassing causal ML, domain adaptation, uncertainty quantification, and multi-objective optimization, illustrating the path forward for harmonizing ML and public sector objectives.
@@ 3 Joint Program in Survey Methodology, University of Maryland, USA
4 Dept. of Software and Information Systems Engineering and Dept. of Epidemiology, Biostatistics and Community Health Sciences, Ben Gurion University of the Negev, Israel
@@ 1 Introduction
Automated decision-making (ADM) systems are increasingly being adopted across the public sector (Algorithm Watch, , 2020; Mitchell et al., , 2021; Levy et al., , 2021), often relying on machine learning (ML) models to address a wide array of problem domains, including critical areas such as predictive policing (Lum and Isaac, , 2016), criminal justice (Angwin et al., , 2016; McKay, , 2020), fraud detection in government (Engstrom et al., , 2020), child abuse prevention (Chouldechova et al., , 2018), tax audit selection (Black et al., , 2022), early warning systems in public schools (Perdomo et al., , 2023), credit scoring (Kozodoi et al., , 2022), profiling of jobseekers (Desiere and Struyven, , 2021; Körtner and Bonoli, , 2023; Bach et al., , 2023), development aid (Kuzmanovic et al., , 2024) and public health (Potash et al., , 2015). Despite expectations of enhancing decision-making by improving reliability, objectivity, efficiency and uncovering factors that traditional institutional processes may overlook, ADM systems face considerable challenges (Barocas et al., , 2019; Coston et al., , 2023; Engstrom et al., , 2020; Wang et al., , 2023; Levy et al., , 2021). Real-world examples frequently demonstrate shortcomings, ranging from racial and gender bias to systems exhibiting poor predictive accuracy leading to flawed decision-making (Obermeyer et al., , 2019; Allhutter et al., , 2020; Angwin et al., , 2016; Dressel and Farid, , 2018; Mayer et al., , 2020). Such unintended consequences are particularly concerning due to their significant impact on individuals’ lives and the potential reinforcement of systemic biases. Recent legislation, such as the European Union’s AI Act, highlights these concerns by establishing regulations for high-risk AI systems (Laux et al., , 2023).
@@
A growing body of literature explores the challenges and potential benefits of employing AI systems to enhance decision-making within the public sector (Sun and Medaglia, , 2019; Zuiderwijk et al., , 2021; Pencheva et al., , 2020; Wirtz et al., , 2019). Moreover, several reviews examine the adoption of AI in government (Levy et al., , 2021), including US federal institutions (Engstrom et al., , 2020) and the EU public sector (van Noordt and Misuraca, , 2022; Algorithm Watch, , 2019). While these reviews cover a wide range of challenges, their focus is on institutional and social implications rather than technical challenges and methodological developments.
@@
One central challenge in implementing algorithmic systems in the public sector is the tension between complex, ambiguous policy objectives and the explicit formalization requirements demanded by ML models (Levy et al., , 2021; Mitchell et al., , 2021; Amarasinghe et al., , 2023; Passi and Barocas, , 2019). Formulating policy objectives and decision-making invariably involves various stakeholders, frequently entangled in a political process characterized by conflicting goals (Levy et al., , 2021; Coyle and Weller, , 2020). Building a system that aligns with the originally intended outcome, defined by complex and ambiguous political compromises (Coyle and Weller, , 2020), necessitates careful consideration of various technical choices during development, including an examination of data assumptions and the model’s integration into the decision-making process. Efforts to address these challenges from a technical perspective are ongoing and focus on connecting methodological ML research with the unique demands of high-stakes decision-making. These efforts explore various subdimensions of this complex issue, including explainability (Amarasinghe et al., , 2023), fairness (Mitchell et al., , 2021; Körtner and Bach, , 2023; Barocas et al., , 2019), target variable bias (Guerdan et al., 2023b, ) and uncertainty (Kaiser et al., , 2022). Furthermore, active research develops frameworks to examine the conditions under which the usage of predictive algorithms for high-stakes decision-making can be justified (Coston et al., , 2023; Wang et al., , 2023).
Unlike many other ML papers focused on decision-making, our paper is exclusively concerned with the issues that emerge when conducting high-stakes decision-making in the public sector. We will discuss ML decision making in a way that is adapted to this specific context by giving a detailed discussion of resource allocation problems commonly encountered in the public sector. These scenarios usually involve a set of individuals, which may refer to individual people, organizations or cases (Kuppler et al., , 2022). For each individual a decision must be made regarding an appropriate intervention from a set of possible interventions. Given that many interventions involve some form of cost and that resources in the public sector are typically scarce, the decision-making process typically revolves around determining whether an individual should receive a particular good or service. For example, we may wish to distribute limited ICU beds among patients, or identify the financial reports that warrant investigation for potential fraud. ML models are then employed to either directly estimate the optimal allocation of interventions, or to predict the individual outcomes under different interventions to inform the downstream decision-making process.
This paper contributes to bridging the gap between ML and decision-making in the public sector by providing a concise yet holistic overview of central methodological decision-points. We present three modeling frameworks and illustrate their use and alignment with decision-making scenarios in the public sector: risk prediction, counterfactual prediction and policy learning. We map each modeling framework to key points at which disjunctions between policy goals and models can arise, and highlight methodological advancements from the technical literature to help avoid common pitfalls in ADM practice. It offers a critical yet productive addition to this field by not only giving a systematic overview of central inflection points in the model-building process but also directly connecting them with promising methodological developments. This paper should thus be of use to any ML practitioner and policymaker seeking an in-depth discussion of the fundamental challenges of building models for public sector decision-making and also serves as a well-structured research guide to the interdisciplinary array of methods available to address these challenges.
@@
We will first explore central technical challenges that occur along the ML pipeline when developing and deploying ML systems to support decision-making in the public sector (Section 2). Second, we will highlight recent methodological developments that exceed the classical supervised ML paradigm, showing promise in addressing the challenges identified (Section 3). In the discussion (Section 4), we place special emphasis on selecting an appropriate modeling approach and target estimand to effectively inform decision-making in a given context.
@@ 2 Defining the Gap: Central Challenges in Connecting ML and Decision-Making
Predictive models for ADM systems are designed to inform decisions in (interaction with) dynamic social contexts, which gives rise to a list of fundamental challenges. This includes questions related to choosing adequate model input, as the effectiveness of any ML model is fundamentally linked to the quality of its training data. Ensuring that this data accurately represents the target population is key to avoid biased and unreliable predictions (Gruber et al., , 2023).
Securing representative data in the public sector, however, is a complex task. Public sector data is incredibly diverse (Dwivedi et al., , 2021; Janssen et al., , 2017), comes in a variety of formats, often lacks structure and encompasses a wide array of data modalities (Dwivedi et al., , 2021). Despite the abundance of data in theory, high-quality data suitable for ML is often not easily available in the public sector (Alexopoulos et al., , 2019; Sun and Medaglia, , 2019). In many countries, the lack of robust infrastructure to enable data sharing and integration of various data sources can hinder the development of ML models (Sun and Medaglia, , 2019; Wirtz et al., , 2019), stemming from issues such as resource constraints, data protection, safety concerns and institutional pushback. These issues are especially concerning for ADM systems, since building a model that is capable of informing future decision-making places significant demands on the training data (Hüllermeier, , 2021; Coston et al., , 2023).
@@
In the following subsections, we discuss key technical challenges associated with both data input and model output that are especially relevant for high-stakes decision-making in the public sector (see Figure 1). These challenges include considering potential distribution shifts between the model’s training and deployment context (Section 2.1), dealing with proxy variables in complex policy settings (Section 2.2) and discerning the impact of past decision-making on the data (Section 2.3). We will also discuss the difficulty of handling multiple potentially conflicting goals (Section 2.4), and the role of human decision-makers whose judgment can potentially overrule the recommendations made by an algorithm (Section 2.5).
@@ 2.1 Distribution Shifts
A fundamental challenge when deploying ML models is that the data on which the system is trained and evaluated may not accurately represent the population of interest in the deployment environment (Kouw and Loog, , 2018; Gruber et al., , 2023). When a model is trained using source data $\mathcal{D}_{\mathcal{S}}$ that is differently distributed than the target data $\mathcal{D}_{\mathcal{T}}$ it will be used on in the relevant application, the model is likely to experience a decline in performance. This phenomenon is commonly referred to as distribution or dataset shift between the joint distribution of the source data $p_{\mathcal{S}}(X,Y)$ and the target distribution $p_{\mathcal{T}}(X,Y)$ of the deployment data. Commonly investigated types of shifts follow from possible decompositions of the joint distribution $p(X,Y)=p(Y|X)p(X)=p(X|Y)p(Y)$, assuming that one component remains invariant, while the other differs (Kouw and Loog, , 2018; Moreno-Torres et al., , 2012). For example, shifts can occur in the covariates $p_{\mathcal{S}}(X)\neq p_{\mathcal{T}}(X)$ or in labels $p_{\mathcal{S}}(Y)\neq p_{\mathcal{T}}(Y)$, with the respective conditional distributions $p(Y|X)$ or $p(X|Y)$ remaining unchanged. Meanwhile, concept shift refers to a change in the relationship between covariates and labels, while keeping prior distributions $p(X)$ or $p(Y)$ constant.
@@
Distribution shifts often result from a biased selection of training data (Moreno-Torres et al., , 2012). For example, it may be more costly to collect relevant data for hard to reach subgroups in the population, leading to them being underrepresented in the data used for training. Selection bias may be introduced through a variety of other mechanisms, for example if past decision policies have led to only certain subgroups receiving an intervention, it will be difficult to assess the interventions’ impact for other individuals. Similarly, deploying a model in a new problem context, for instance in a different geographic region, may also lead to a distribution shift.
Medical prediction models are instructive here, as population health and healthcare practices change dramatically over time and place, making these models particularly susceptible to distribution shifts. For example, the well-known Framingham models used for cardiovascular disease risk prediction were found to overestimate risk within the European population (Damen et al., , 2019). This overestimation may be attributable to advances in available preventive treatments since the models were initially developed using data from the second half of the 20th century (Damen et al., , 2019).
@@
However, even a perfect selection of training data does not guarantee long-term robustness. As changes in the deployment environment occur, the initially accurate data may become increasingly outdated, likely causing the performance of a model to degrade over time (Moreno-Torres et al., , 2012). When a model is used to inform future decision-making, its continued deployment may itself be a source of distribution shift. For instance, individuals might strategically manipulate attributes that are not causally related to the true outcome but are correlated to improve predictions in their favor, often worsening the model’s accuracy in predicting the true outcome of interest (Hardt et al., 2016a, ). This is a known challenge when making use of models to support enforcement decisions in government, such as financial fraud detection. In order to evade detection, certain regulatory subjects will adapt to a given system, thereby requiring continuous updates to maintain effectiveness (Engstrom et al., , 2020). The performance of a model may also be impacted by more sudden changes in the deployment environment. The introduction of a new policy can influence how new training data is collected and labeled, and unexpected events, such as the COVID-19 pandemic (Singh et al., , 2021), can reduce the prediction accuracy of a model.

### v3

# Bridging the Gap: Towards an Expanded Toolkit for AI-Driven Decision-Making in the Public Sector

@@ Abstract
AI-driven decision-making systems are becoming instrumental in the public sector, with applications spanning areas like criminal justice, social welfare, financial fraud detection, and public health. While these systems offer great potential benefits to institutional decision-making processes, such as improved efficiency and reliability, these systems face the challenge of aligning machine learning (ML) models with the complex realities of public sector decision-making. In this paper, we examine five key challenges where misalignment can occur, including distribution shifts, label bias, the influence of past decision-making on the data side, as well as competing objectives and human-in-the-loop on the model output side. Our findings suggest that standard ML methods often rely on assumptions that do not fully account for these complexities, potentially leading to unreliable and harmful predictions. To address this, we propose a shift in modeling efforts from focusing solely on predictive accuracy to improving decision-making outcomes. We offer guidance for selecting appropriate modeling frameworks, including counterfactual prediction and policy learning, by considering how the model estimand connects to the decision-maker’s utility. Additionally, we outline technical methods that address specific challenges within each modeling approach. Finally, we argue for the importance of external input from domain experts and stakeholders to ensure that model assumptions and design choices align with real-world policy objectives, taking a step towards harmonizing AI and public sector objectives.
@@ 3 Joint Program in Survey Methodology, University of Maryland, USA
4 Dept. of Software and Information Systems Engineering and Dept. of Epidemiology, Biostatistics and Community Health Sciences, Ben-Gurion University of the Negev, Israel
Highlights:
  • Machine learning (ML) is frequently used to support decision-making in the public sector
  • A key challenge is the misalignment between ML models and the realities of public sector decision-making
  • We analyze five challenges to investigate how misaligned technical assumptions can lead to erroneous decision-making
  • We argue for a shift from focusing solely on predictive accuracy to improving decision-making outcomes
  • We offer guidance on selecting the right modeling framework, with a focus on causal machine learning and stakeholder input
@@ 1 Introduction
Automated decision-making (ADM) systems are increasingly being adopted across the public sector (Algorithm Watch, , 2020; Mitchell et al., , 2021; Levy et al., , 2021), often relying on AI models to address a wide array of problem domains, including critical areas such as predictive policing (Lum and Isaac, , 2016), criminal justice (Angwin et al., , 2016; McKay, , 2020), fraud detection in government (Engstrom et al., , 2020), child abuse prevention (Chouldechova et al., , 2018), tax audit selection (Black et al., , 2022), early warning systems in public schools (Perdomo et al., , 2023), credit scoring (Kozodoi et al., , 2022), profiling of job seekers (Desiere and Struyven, , 2021; Körtner and Bonoli, , 2023; Bach et al., , 2023), development aid (Kuzmanovic et al., , 2024) and public health (Potash et al., , 2015). Despite expectations of enhancing decision-making by improving reliability, objectivity, efficiency and uncovering factors that traditional institutional processes may overlook, ADM systems face considerable challenges (Barocas et al., , 2023; Coston et al., , 2023; Engstrom et al., , 2020; Wang et al., , 2023; Levy et al., , 2021). Real-world examples demonstrate shortcomings, ranging from racial and gender bias to systems exhibiting poor predictive accuracy leading to flawed decision-making (Obermeyer et al., , 2019; Allhutter et al., , 2020; Angwin et al., , 2016; Dressel and Farid, , 2018; Mayer et al., , 2020). Such unintended consequences are particularly concerning due to their significant impact on individuals’ lives and the potential reinforcement of systemic biases. Recent legislation, such as the European Union’s AI Act, highlights these concerns by establishing regulations for high-risk AI systems (Laux et al., , 2023).
@@
A growing body of literature explores the challenges and potential benefits of employing AI systems to enhance decision-making within the public sector (Sun and Medaglia, , 2019; Zuiderwijk et al., , 2021; Pencheva et al., , 2020; Wirtz et al., , 2019). Moreover, several reviews examine the adoption of AI in government (Levy et al., , 2021), including US federal institutions (Engstrom et al., , 2020) and the EU public sector (van Noordt and Misuraca, , 2022; Algorithm Watch, , 2019). These reviews cover a wide range of challenges, primarily focusing on institutional, ethical and legal implications of using ADM in the public sector.
In this article, we focus on challenges that arise from a misalignment between the technical assumptions underlying machine learning (ML) models and the realities of decision-making in complex public sector environments. Specifically, we will discuss AI-driven decision-making used for the allocation of scarce resources in the public sector, where decisions involve determining whether individuals qualify to receive specific interventions or services (Kuppler et al., , 2022). Our focus is on ADM systems that do not rely on manually encoded rules, but rather use supervised ML models to learn patterns from historical data to predict relevant outcomes that inform decision-making. Although ML approaches can vary widely, ranging from support vector machines to neural networks, we aim to keep our discussion relevant across different models by exploring the general limitations and challenges of using supervised ML for public sector decision-making. Throughout the text, we use terms like AI, ML and predictive algorithm interchangeably to refer to the computational model underlying the ADM system.
Decision-making in these environments often takes place in dynamic, evolving social contexts, which can conflict with the explicit formalization requirements demanded by ML models (Levy et al., , 2021; Mitchell et al., , 2021; Amarasinghe et al., , 2023; Passi and Barocas, , 2019). Technical choices made during model development often rest on implicit assumptions, such as stable data distributions and a straightforward link between prediction and decision-making, that may not hold in these complex settings. For example, policy objectives are often shaped by multiple stakeholders, political compromises and competing goals (Levy et al., , 2021; Coyle and Weller, , 2020), making it difficult to translate them into clearly defined objectives for ML systems. When the assumptions behind the technical model construction do not align with the deployment context, there is a risk of developing systems that fail to capture the complexities of real-world decision-making, potentially leading to adverse outcomes upon model deployment.
Consider, for example, a public employment service (PES) office that aims to determine which job seekers should participate in job programs to increase their re-integration chances. The PES wants to deploy ML to learn the optimal assignment of support to job seekers based on data collected as part of their daily operations. However, the PES now faces two critical sets of interconnected complications: first, while their data may include detailed records of labor market histories, they are operating in a complex and dynamic social environment which raises questions of distribution shift, feedback loops and the challenge of accounting for the effect of competing (current) and previous job support programs. Second, the PES needs to ensure that the predictions can effectively be integrated in their current decision-making practices. This may require model guarantees to build caseworker trust in the predictions and ensuring that other relevant objectives and constraints are sufficiently incorporated in the system. All these issues require careful consideration in the model design choices. A misalignment between technical assumptions and problem setting, such as building a model under the implicit assumption that labor market characteristics remain invariant, may result in unintended consequences, such as an allocation mechanism that might become unreliable over time.
@@
Efforts to analyze challenges from a technical perspective are ongoing and focus on connecting methodological AI research with the unique demands of high-stakes decision-making. These efforts explore various subdimensions of this complex issue, including training data quality (Shahbazi et al., , 2023), target variable bias (Guerdan et al., 2023b, ) and uncertainty (Gruber et al., , 2023; Kaiser et al., , 2022). Furthermore, active research develops frameworks to examine the conditions under which the usage of predictive algorithms for high-stakes decision-making can be justified (Coston et al., , 2023; Wang et al., , 2023).
In this work, we identify and analyze misalignments that commonly occur between ML models and public sector decision-making. Guided by the ‘ADM process model’ (Gerdon et al., , 2022), we focus on how models connect with their wider real-world deployment context by examining both the data assumptions (model input) and how models are integrated into the decision-making process (model output). Using the lens of misalignment developed here, we build on the recent technical literature on ML and decision-making to isolate five specific challenges that we consider to exemplify the type of issues that can occur at these two interfaces: distribution shift, label bias and the influence of past decision-making on the input side, and competing objectives and constraints and human-in-the-loop interactions on the output side. We analyze each of these challenges to better understand how misaligned technical assumptions can lead to erroneous decision-making and adverse outcomes for affected individuals in public sector environments.
Through our analysis, we find that standard ML methods often rely on assumptions that do not fully account for the complexities of public sector decision-making. In response, we propose a shift in modeling efforts from focusing solely on predictive accuracy to improving decision-making outcomes. We argue that achieving this shift may, in certain cases, require alternative modeling techniques that extend purely predictive models, and more directly center on the goal of decision-making. With this in mind, we highlight promising developments in causal machine learning, including counterfactual prediction and policy learning. Within each modeling framework, we summarize technical methods that provide (partial) solutions to the identified challenges. To guide practitioners in selecting the right approach, we clarify the assumptions underlying each framework, specifically addressing, how the model estimand connects to the utility of the decision-maker and the data and assumptions required for reliable estimation.
By examining these frameworks through the lens of public sector decision-making, we want to encourage technical practitioners to carefully consider the assumptions behind different modeling approaches and expand their toolbox to include methods that may be better suited for complex, real-world decision-making. For policymakers, domain experts and other stakeholders, we outline which external input is important to help model developers make the right assumptions to inform model design.
While we do discuss several risks resulting from the assumptions made during model development, zooming out to the institutional and societal context raises more complex issues. For instance, institutional and cultural biases embedded in historical data as well as data collection methods and processing can significantly contribute to discriminatory decision-making (Janssen and Kuk, , 2016; Fountain, , 2022). Algorithmic systems may also reinforce existing structural inequalities by formalizing problematic decision-making practices (Kolkman, , 2020) or empowering institutions with unjust goals. The continued digitization of bureaucratic processes, particularly when multiple institutions and systems interact, can create new risks, such as making it harder to correct errors across systems or systematically excluding specific user groups (Peeters and Widlak, , 2018). However, we consider addressing misalignments between assumptions made during model development and the deployment context to be essential for avoiding harmful model design, making it a necessary (though not sufficient) condition for the fair development of AI-driven decision making in the public sector.
@@
This paper is structured as follows. We first explore central (mis)alignment challenges that occur along the ML pipeline when developing and deploying AI systems to support decision-making in the public sector (Section 2). Second, we highlight recent methodological developments that exceed the classical supervised ML paradigm, showing promise in addressing the challenges identified (Section 3). Third, we discuss the selection of an appropriate modeling approach in a given deployment context (Section 4). In the discussion, we address broader issues related to ADM in the public sector, specifically highlighting the importance of domain expertise and stakeholder input (Section 5). Finally, Section 6 provides a concise summary of our findings.
@@ 2 Defining the Gap: Central Challenges in Connecting ML and Decision-Making
Predictive models for ADM systems are designed to inform decisions in (interaction with) dynamic social contexts, which gives rise to a list of fundamental challenges. This includes questions related to choosing adequate model input, as the effectiveness of any ML model is fundamentally linked to the quality of its training data. Ensuring that this data accurately represents the target population is key to avoid biased and unreliable predictions (Gruber et al., , 2023). Securing representative data in the public sector, however, is a complex task. Public sector data is incredibly diverse (Dwivedi et al., , 2021; Janssen et al., , 2017), comes in a variety of formats, often lacks structure and encompasses a wide array of data modalities (Dwivedi et al., , 2021). Despite the abundance of data in theory, high-quality data suitable for ML is often not easily available in the public sector (Alexopoulos et al., , 2019; Sun and Medaglia, , 2019). In many countries, the lack of robust infrastructure to enable data sharing and integration of various data sources can hinder the development of ML models (Sun and Medaglia, , 2019; Wirtz et al., , 2019), stemming from issues such as resource constraints, data protection, safety concerns and institutional pushback. These issues are especially concerning for ADM systems, since building a model that is capable of informing future decision-making places significant demands on the training data (Hüllermeier, , 2021; Coston et al., , 2023).
@@
In the following subsections, we discuss key challenges associated with both data input and model output that are especially relevant for high-stakes decision-making in the public sector (see Figure 1). These challenges include considering potential distribution shifts between the model’s training and deployment context (Section 2.1), dealing with proxy variables in complex policy settings (Section 2.2) and discerning the impact of past decision-making on the data (Section 2.3). We will also discuss the difficulty of handling multiple potentially conflicting goals (Section 2.4), and the role of human decision-makers whose judgment can potentially overrule the recommendations made by an algorithm (Section 2.5).
@@ 2.1 Distribution Shifts
A key challenge when using ML models is to ensure that they perform well in real-world situations. Often, the data used to train and evaluate the model will not fully represent the actual population in the environment where the model will be deployed (Kouw and Loog, , 2018; Gruber et al., , 2023). This mismatch between the distribution of training and deployment data is commonly referred to as distribution shift, and can lead to a significant overestimation of a model’s performance (Kouw and Loog, , 2018; Gruber et al., , 2023). In other words, the model may learn patterns in the training data that do not generalize well to the deployment data, causing it to perform poorly in practice. Distribution shifts are especially challenging in the public sector, where sourcing reliable data can be particularly difficult. Models are often deployed in complex and evolving social contexts, and limited resources, such as a shortage of technical staff (Wirtz et al., , 2019), make it difficult for public institutions to monitor model performance for unexpected distribution shifts.
There are several types of shifts that can occur (Kouw and Loog, , 2018; Moreno-Torres et al., , 2012). For example, the distribution of input covariates, such as age, income or educational background, might vary if a model is trained in one geographic region but deployed in another. An even harder type of shift to address is when the relationship between input covariates and the outcome changes, making the model’s predictions less applicable in the new setting.
@@
Distribution shifts often result from a biased selection of training data (Moreno-Torres et al., , 2012). For example, it may be more costly to collect relevant data for hard to reach subgroups in the population (Tourangeau, , 2019), leading to them being underrepresented in the data used for training. Selection bias may be introduced through a variety of other mechanisms, for example if past decision policies have led to only certain subgroups receiving an intervention, it will be difficult to assess the interventions’ impact for other individuals.
@@
Even a comprehensive selection of training data does not guarantee long-term robustness. As changes in the deployment environment occur, the initially accurate data may become increasingly outdated, likely causing the performance of a model to degrade over time (Moreno-Torres et al., , 2012). For instance, labor market characteristics might change over time, making a model trained on older data for predicting unemployment less accurate. When a model is used to inform future decision-making, its continued deployment may itself be a source of distribution shift. For instance, individuals might strategically manipulate attributes that are not causally related to the true outcome but are correlated to improve predictions in their favor, often worsening the model’s accuracy in predicting the true outcome of interest (Hardt et al., 2016a, ). This is a known challenge when making use of models to support enforcement decisions in government, such as financial fraud detection. In order to evade detection, certain regulatory subjects will adapt to a given system, thereby requiring continuous updates to maintain effectiveness (Engstrom et al., , 2020). The performance of a model may also be impacted by more sudden changes in the deployment environment. The introduction of a new policy can influence how new training data is collected and labeled, and unexpected events, such as the COVID-19 pandemic (Singh et al., , 2021), can reduce the prediction accuracy of a model.